OpenStudy (anonymous):

What is the simplified form of the expression the square root of the quantity 4 times c to the sixteenth power. - the square root of c squared. + 6the square root of c squared. + the square root of the quantity 9 times c squared.?

OpenStudy (campbell_st):

\[\sqrt{4c^{16}} - \sqrt{c^2} + 6\sqrt{c} + 9\sqrt{c^2}\] simplify the radicals \[\sqrt{4c^{16}} = 2c^8\] \[\sqrt{c^2} = c\] \[9\sqrt{c^2} = 9c\] make the substitution and simplify

OpenStudy (anonymous):

can you explain further please?

OpenStudy (campbell_st):

ok so after substituting \[2c - c + 6\sqrt{c} + 9c = (2c - c + 9c) + 6\sqrt{c}\] just simplify the part in brackets

OpenStudy (anonymous):

2c+9c

OpenStudy (anonymous):

right?